A bias-neutral venture capitalist is contemplating investing $6 million in a start-up company. Company net income at end of year six is expected to be $5 million. Given the six year length of his investment, his firm requires a 80% internal rate of return compounded annually. Given that the market for the company’s product line is expected to be biotech and have multiple uses, the price to earnings ratio is expected to be 45. Neither future investments nor change in earnings projection is expected prior to the end of the investment term. There are currently 5 million shares outstanding.
Answer #1

The VC must realise a 80% return on investment after 6 years.

Therefore,

His investment value after 6 years = Investment*(1+r)^{n}

Investment worth after 6 years = 6*(1+0.8)^{6} = $204.0733 Mn
Net Income(year 6) = $5Mn

Expected Price/ Earning = 45

Value of company at end of year 6= Net Income * P/E ratio = 45*$5Mn = $225Mn

For VC to receive $204.0733Mn at end of year 6 out of $225Mn Company value,

Ownership% required = 204.0733/225 = 90.69%